# Superradiant Axionic Black-Hole Clouds as Seeds for Graviton Squeezing

### v1

@@ Abstract
It is shown that both, standard general relativity (GR) and Chern-Simons (CS) gravity, the latter containing chiral gravitational anomaly terms, seed the production of pairs of entangled gravitons in a multi-mode squeezed state. This involves the interaction of gravitons with the axionic cloud surrounding a superradiant Kerr (rotating) black hole background. The order of magnitude of the squeezing effect, specifically the number of graviton excitations in the squeezed vacuum, is estimated in the non-relativistic limit, relevant for the superradiance process. It is found analytically that the squeezing from the GR process of annihilation of two axions into two gravitons, dominates, by many orders of magnitude, that coming from the axion decay into two gravitons, induced by the higher-derivative CS term. It is also shown that significant squeezing effects are produced in the case of long-lived axionic clouds, whose lifetimes are much longer than the timescale for which superradiance is effective.
@@
Introduction. It is well known that a mathematically and physically consistent theory combining both quantum theory and general relativity still eludes us. However, there is no consensus on whether the gravitational field is quantized just like the other fundamental fields of nature [1, 2, 3, 4, 5, 6, 7, 8, 9]. In contrast to the detection of a single photon, it is a near-impossible task [10, 11] to detect a single graviton. However, collective states of gravitons might be detectable. GR can be consistently treated as an effective field theory (EFT) [12, 13], within the framework of the background-field method [14, 15]. The validity of the EFT requires energies below the ultraviolet (UV) cut-off, whose upper limit is given by the (reduced) Planck mass scale, $M_{\rm Pl}=\kappa^{-1}=2.435\times 10^{18}$ GeV. This defines the perturbative quantum gravity (PQG) approach, in which one considers a fixed classical background and small quantum perturbations about it. In the PQG regime, non-renormalizabilty issues are avoided and one can make potentially observable predictions. These include properties of gravitational waves (GWs), which were observed for the first time in 2015 [16]. The future prospects for optimizing both Earth-based and space-based GW detectors promise a new era of GW astronomy [17]. A new cross-fertilization with ideas from quantum optics [18, 19, 20, 21, 22, 23] may lead to the detection of the quantum nature of gravity in GWs.
In future GW interferometric detectors [24, 25], quantum fluctuations of the gravitational field might be detectable as an appropriate kind of quantum noise, depending on the state of gravitons. Squeezed states, which are states with no classical analogue, may lead to detectable signatures for a large enough squeezing parameter. For example , the non-classicality in GWs can be explored by employing a Hanbury-Brown-Twiss (HBT) interferometer [26, 27], which can reveal sub-Poissonian statistics for the number of gravitons in squeezed coherent states. However HBT analysis is purely theoretical, with no experimental implementation since it depends on graviton counting. Another approach uses the electromagnetic field cavity as an optomechanical GW detector [28, 29], where quantum properties of the graviton states can be revealed by statistical measurements. For a single-mode squeezed vacuum state to have an observational signature, a squeezing parameter [21] $r\sim 78$ is required, corresponding to an excited vacuum containing $\langle N\rangle=\sinh^{2}r\sim 10^{67}$ gravitons [28].
@@
The axionic fields in the BH cloud can have diverse origins. Our axions are either model-independent string-axions (Kalb-Ramond (KR) axions)  [45, 43] or are equivalent to the totally antisymmetric components of torsion in Einstein-Cartan models [43, 46] for contorted spacetime geometries. The squeezing effect is proportional to the coupling strength of the axion to the (gravitational) CS anomaly.
Superradiance. It is known that a rotating black hole (BH) has a superradiant instability, where bosonic particles are abundantly produced [40]. Upon extracting rotational energy from the BH, these particles exist in quasibound states around the BH interpreted as a condensate-like distribution. In a background of a Kerr BH [47] of mass $\mathcal{M}$, angular momentum $\mathcal{J}_{H}$ and spin parameter $\alpha=\mathcal{J}_{H}/\mathcal{M}$, the Klein-Gordon equation (KG) for a massive (pseudo)scalar field, of mass $\mu_{b}$, admits quasibound states, which are labeled by integer numbers $\left(n,l,m\right)$. We shall denote as $r_{\pm}$ the outer ($+$) and inner (--) event horizons, respectively. We shall work in the non-relativistic regime [48], that is, when the Compton wavelength of the axion $\lambda_{C}\sim 1/\mu_{b}$ is much larger than the black hole radius and the BH-axion binding energy is small compared to the axion rest mass. In this regime, the angular part of the solution of KG is described by spherical harmonics, while the effective radial equation reduces to a Coulomb-like problem, with solutions resembling the wavefunctions of the hydrogen atom [49]. The imaginary part of the frequency $\omega_{I}$ drives the superradiant instability with a growth $\sim e^{\omega_{I}t}$, provided the real part of the frequency, $\omega$, satisfies the superradiant condition [50] $\omega<m\Omega_{H}$, where $\Omega_{H}$ is the angular velocity of the event horizon. For the real part of the frequency, we have [40]:
@@
[…]
where $N_{nlm}$ is chosen to be the number of axions in the respective state, with $\Psi_{nlm}(\vec{x})$ obeying the normalization $\int d^{3}x|\Psi_{nlm}|^{2}=1$. As long as the superradiance condition $\omega<m\Omega_{H}$ is satisfied, the axionic-cloud will grow at a rate faster than the evolution timescale of the BH. The most dominant mode corresponds to the “$2p$-axion state” $(n=2,\ l=m=1)$. Regarding timescales relevant to superradiance, there are three main processes that govern the evolution: the superradiant growth of the instability, the accretion of black holes and GW emission [51, 52]. The instability time $\tau_{s}$ of superradiance for the $2p$-state scales as [48]:
@@
where we consider highly rotating BHs, $\alpha/(G\mathcal{M})\sim\mathcal{O}(1)$ and $a_{\mu}=0.1$, consistent with the non-relativistic approximation [40, 48]. Note that $\tau_{s}$ is a separate timescale from the oscillation period of the cloud $\tau\sim\mu_{b}^{-1}$, and therefore, a quasi-stationary approximation is valid. The evolution of the superradiant instability unfolds in two stages: first, the axionic condensate evolves due to superradiance around the BH background, on a timescale given by $\tau_{s}$ (2). Then, the cloud emits energy via GWs, a process characterized by the relevant timescale $\tau_{GW}$. There is an important separation of these two scales ($\tau_{GW}\gg\tau_{s}$), ensuring two basic physical features; i) the cloud can evolve to saturation without being depleted through GWs emission and ii) the axionic-condensate is very long lived, as supported by various numerical simulations [51, 52, 53, 54]. Since the lifetime $T$ of the cloud has to take into account subprocesses with widely separated time scales, it is considered to be a phenomenological parameter.
@@
where $A$ denotes the coupling constant, and $\widetilde{R}_{\mu\nu\rho\sigma}=\frac{1}{2}R_{\mu\nu\alpha\beta}\ \epsilon^{\alpha\beta}\!_{\rho\sigma}$ is the dual of the Riemann tensor, with $\epsilon_{\mu\nu\rho\sigma}$ the covariant Levi-Civita tensor.
@@
where $\hat{\alpha}_{\lambda,\vec{k}},\hat{\alpha}^{\dagger}_{\lambda,\vec{k}}$ denotes the dimensionless annihilation/creation operators obeying the usual commutation relations, $[\hat{\alpha}_{\lambda,\vec{k}},\hat{\alpha}^{\dagger}_{\lambda^{\prime},\vec{k}^{\prime}}]=\delta_{\lambda\lambda^{\prime}}\delta_{\vec{k}\vec{k}^{\prime}}$, where $\lambda=L,R$ denotes left (L) and right (R) polarization respectively, $f_{k}$ is the (dimensionless) single graviton strain, $f_{k}\equiv\kappa(2V\Omega_{k})^{-\frac{1}{2}}$, while $\Omega_{k}\equiv k$ is the frequency of GW with momentum $\vec{k}$. The polarization tensors obey the normalization $\epsilon^{*(\lambda)}_{ij}\epsilon_{(\lambda^{\prime})}^{ij}=2\delta_{\lambda\lambda^{\prime}}$. The interaction linear in the graviton (5) is responsible for the production of graviton coherent states [61]. The remaining interactions correspond to processes including two gravitons; the first one (6), stemming from GR, describes two axions annihilating into two gravitons, while the second one (Superradiant Axionic Black-Hole Clouds as Seeds for Graviton Squeezing), associated with the anomaly, corresponds to an axion decay into two gravitons.
@@
$\hat{H}_{\text{int}}=\sum_{I,J}e^{i(\Omega_{k}+\Omega_{k^{\prime}}-E)t}\mathcal{F}_{IJ}\hat{\alpha}^{\dagger}_{I}\hat{\alpha}^{\dagger}_{J}+h.c+\dots{}$ UNKNOWN (9)
@@
$I_{ij}(\vec{k},\vec{k}^{\prime})=\int d^{3}\vec{x}\ \partial_{i}\Psi_{2p}(\vec{x})\partial_{j}\Psi_{2p}(\vec{x})e^{-i(\vec{k}+\vec{k}^{\prime})\cdot\vec{x}}\,.$ (14)
derived from the source. Since two axions are involved in the process, $E=2\mu_{b}$. The axion occupation number $N_{2p}$ acts as an enhancement parameter (in analogy to the quantum-optics SFWM). The angular and polarization correlations of the gravitons emitted from the axionic cloud are determined from (13). Considering gravitons emitted in the $y-z$ plane with $\vec{k}+\vec{k}^{\prime}$ corresponding to angles $\theta_{k+k^{\prime}}=\varphi_{k+k^{\prime}}=\pi/2$, we can express (13) in terms of the angle $\Delta\theta$ between the two gravitons (c.f. figure 2).
Figure 2: Angular and polarization correlations for the GR interaction. The $2p$-state results in the asymmetry between the LL and RR pairs. The plot corresponds to $a_{\mu}=0.1$.
@@
×$\times\left([e^{(3)}(\vec{k}^{\prime})]_{m}\ e^{(\lambda)}_{mj}(\vec{k})\ e^{(\lambda^{\prime})}_{jl}(\vec{k}^{\prime})[e^{(3)}(\vec{k})]_{l}\ +\right.$
@@
Figure 3: Angular and polarization correlations for the CS interaction. Only pairs of opposite polarizations are produced; Maximal entanglement occurs between the L and R polarizations. The plot corresponds to $a_{\mu}=0.1$.
@@
×$\times\sum_{\lambda,\lambda^{\prime}}\int d\tilde{k}d\tilde{k}^{\prime}\ \tilde{k}\tilde{k}^{\prime}\delta\left(\tilde{k}+\tilde{k}^{\prime}-2\right)\Big{|}\widetilde{\mathcal{I}}^{(GR)}_{(\lambda,\vec{k})(\lambda^{\prime},\vec{k}^{\prime})}\Big{|}^{2}\ ,$ (21)
where $\tilde{k}^{(\prime)}=k^{(\prime)}/\mu_{b}$ and =~$\mathcal{I}=\widetilde{\mathcal{I}}/\mu_{b}^{2}$. In order to estimate the number of gravitons in the squeezed vacuum state, we consider $\vec{q}=-(\vec{k}+\vec{k}^{\prime})$ lying on $\theta_{k+k^{\prime}}=\varphi_{k+k^{\prime}}=\pi/2$ (c.f. figure 2). The main contribution to the integral (21) comes from momentum vectors $\vec{k}$, $\vec{k}^{\prime}$ satisfying $k\approx k^{\prime}\approx\mu_{b}$, whilst the maximum contribution corresponds to their relative angle $\Delta\theta=0.964\pi$. The integrations over the solid angles will yield a factor at most of order $\left(4\pi\right)^{2}$, implying the following upper limit,
@@
Note here that the large suppression induced by the ratio $\big{(}\mu_{b}/M_{\rm Pl}\big{)}$ is compensated by the number of axions in the cloud , $N_{2p}\left(\mu_{b}/M_{\rm Pl}\right)^{2}\approx 10^{-3}a_{\mu}^{2}=10^{-5}$ (see (3)).
This is not the case for the anomaly induced interaction. In an effective field theory, the CS-coupling to the axion is suppressed below the cut-off scale of the theory. For the string-inspired KR axion coupling [43] $A\sim 10^{-2}M_{\rm Pl}/M^{2}_{\text{s}}$, with $M_{\text{s}}$ the string scale. As in the GR case, we argue (c.f. figure 3):
@@
It is evident from (22) and (23) that the lifetime $T$ of the axionic cloud is also a parameter able to induce significant squeezing. In contrast to the short lived quasi-normal modes (QNMs) [69], the longevity of axionic clouds seems to overcome limitations and produce appreciable squeezing effects. We shall identify $T$ with the lifetime of the cloud, $T=\tau_{\text{cloud}}$. Although the lifetime of the axionic condensate is not explicitly defined in the literature, many numerical simulations have been carried out [51, 52, 53, 70], and all seem to agree that there is a clear separation of scales, i.e. $\tau_{\text{cloud}}\gg\tau_{s}$. For example, in [70], the lifetime of the cloud results in a factor $T\sim 10^{7}\tau_{s}$, for which, on using (2) and  (22), yields,
$\sum_{I,J}|\mathcal{G}^{(GR)}_{IJ}|^{2}\sim\mathcal{O}(10)\ .$ (24)
Hence, the total number of gravitons in the squeezed vacuum state will be exponentially enhanced, resembling the $\langle N_{gr}\rangle=\sinh^{2}r$ behavior of the one-mode case [21]. In contrast, the CS anomaly-induced squeezing (23) is still highly suppressed, compared to that induced by GR, since higher curvature interactions must remain subleading at low energy scales [12], determined by the axion mass, $\mu_{b}$.
Our analysis above has been restricted to the flat-background approximation and crucially depends on the longevity of the axionic clouds. However, the approach can be extended to take into account curvature effects around the BH, which may provide an additional enhancement parameter.

### v2

@@ Abstract
It is shown that both, standard general relativity (GR) and Chern-Simons (CS) gravity, the latter containing chiral gravitational anomaly terms, seed the production of pairs of entangled gravitons in a multi-mode squeezed state. This involves the interaction of gravitons with the axionic cloud surrounding a superradiant Kerr (rotating) black hole background. The order of magnitude of the squeezing effect, specifically the number of graviton excitations in the squeezed vacuum, is estimated in the non-relativistic limit, relevant for the superradiance process. It is found analytically that the squeezing from the GR process of annihilation of two axions into two gravitons, dominates, by many orders of magnitude, that coming from the axion decay into two gravitons, induced by the higher-derivative CS term. It is also shown that significant squeezing effects are produced in the case of long-lived axionic clouds, whose lifetimes are much longer than the timescale for which superradiance is effective. A brief discussion on current exclusion (for the first time) of very-long axion-cloud lifetimes, through comparison of our results with current LIGO data, as well as potential detection of such effects in future intereferometers is also given.
@@
Introduction. It is well known that a mathematically and physically consistent theory combining both quantum theory and general relativity still eludes us. However, there is no consensus on whether the gravitational field is quantized just like the other fundamental fields of nature [1, 2, 3, 4, 5, 6, 7, 8, 9]. In contrast to the detection of a single photon, it is a near-impossible task [10, 11, 12] to detect a single graviton. However, collective states of gravitons might be detectable. GR can be consistently treated as an effective field theory (EFT) [13, 14], within the framework of the background-field method [15, 16]. The validity of the EFT requires energies below the ultraviolet (UV) cut-off, whose upper limit is given by the (reduced) Planck mass scale, $M_{\rm Pl}=\kappa^{-1}=2.435\times 10^{18}$ GeV. This defines the perturbative quantum gravity (PQG) approach, in which one considers a fixed classical background and small quantum perturbations about it. In the PQG regime, non-renormalizabilty issues are avoided and one can make potentially observable predictions. These include properties of gravitational waves (GWs), which were observed for the first time in 2015 [17]. The future prospects for optimizing both Earth-based and space-based GW detectors promise a new era of GW astronomy [18]. A new cross-fertilization with ideas from quantum optics [19, 20, 21, 22, 23, 24] may lead to the detection of the quantum nature of gravity in GWs.
In future GW interferometric detectors [25, 26], quantum fluctuations of the gravitational field might be detectable as an appropriate kind of quantum noise [27], depending on the state of gravitons. Squeezed states, which are states with no classical analogue, may lead to detectable signatures for a large enough squeezing parameter. For example , the non-classicality in GWs can be explored by employing a Hanbury-Brown-Twiss (HBT) interferometer [28, 29], which can reveal sub-Poissonian statistics for the number of gravitons in squeezed coherent states. However HBT analysis is purely theoretical, with no experimental implementation so far, since it relies on delicate intensity-correlation techniques related to indirect graviton counting. Another approach uses the electromagnetic field cavity as an optomechanical GW detector [30, 31], where quantum properties of the graviton states can be revealed by statistical measurements. For a single-mode squeezed vacuum state to have an observational signature, a larger than one squeezing parameter [22] is required. In  [31] the authors considered very large squeezing parameters of order $r\sim 78$, corresponding to an excited vacuum containing $\langle N\rangle=\sinh^{2}r\sim 10^{67}$ gravitons, as examples of potentially detectable graviton squeezing in future GW detectors. Nonetheless, the non-observation of squeezed GW states by the LIGO/Virgo GW interferometer [17, 32] imposes stringent upper bounds on the GW-squeezing parameter $r<41$ [33]. The allowed region of $r$ still leaves plenty of room for sufficiently large squeezing parameters, and, as we shall see, it is compatible with our mutli-mode-squeezed-graviton considerations in this work.
@@
The axionic fields in the BH cloud can have diverse origins. Our axions are either string-compactification axions or string-model-independent axions (Kalb-Ramond (KR) axions)  [49, 47] or are equivalent to the totally antisymmetric components of torsion in Einstein-Cartan models [47, 50] for contorted spacetime geometries, provided mechanisms for making all such axions massive are in place.
Superradiance. It is known that a rotating black hole (BH) has a superradiant instability, where bosonic particles are abundantly produced [44]. Upon extracting rotational energy from the BH, these particles exist in quasibound states around the BH interpreted as a condensate-like distribution. In a background of a Kerr BH [51] of mass $\mathcal{M}$, angular momentum $\mathcal{J}_{H}$ and spin parameter $\alpha=\mathcal{J}_{H}/\mathcal{M}$, the Klein-Gordon equation (KG) for a massive (pseudo)scalar field, of mass $\mu_{b}$, admits quasibound states, which are labeled by integer numbers $\left(n,l,m\right)$. We shall denote as $r_{\pm}$ the outer (++) and inner (-) event horizons, respectively. We shall work in the non-relativistic regime [52], that is, when the Compton wavelength of the axion $\lambda_{C}\sim 1/\mu_{b}$ is much larger than the black hole radius and the BH-axion binding energy is small compared to the axion rest mass. In this regime, the angular part of the solution of KG is described by spherical harmonics, while the effective radial equation reduces to a Coulomb-like problem, with solutions resembling the wavefunctions of the hydrogen atom [53]. The imaginary part of the frequency $\omega_{I}$ drives the superradiant instability with a growth $\sim e^{\omega_{I}t}$, provided the real part of the frequency, $\omega$, satisfies the superradiant condition [54] $\omega<m\Omega_{H}$, where $\Omega_{H}$ is the angular velocity of the event horizon. For the real part of the frequency, we have [44]:
@@
$b(t,r,\theta,\phi)=\sum_{nlm}e^{-i\omega_{n}t}\sqrt{\frac{N_{nlm}}{2\mu_{b}}}\Psi_{nlm}+\text{c.c}\ ,$ (1)
where $N_{nlm}$ is chosen to be the number of axions in the respective state, with $\Psi_{nlm}(\vec{x})$ obeying the normalization $\int d^{3}x|\Psi_{nlm}|^{2}=1$. As long as the superradiance condition $\omega<m\Omega_{H}$ is satisfied, the axionic-cloud will grow at a rate faster than the evolution timescale of the BH (see discussion below). The most dominant mode corresponds to the “$2p$-axion state” (n=2,l=m=1)(n=2,\ l=m=1). Regarding timescales relevant to superradiance, there are three main processes that govern the evolution: the superradiant growth of the instability, the accretion of black holes and GW emission [55, 56]. The instability time $\tau_{s}$ of superradiance for the $2p$-state scales as [52]:
@@
where we consider highly rotating BHs, $\alpha/(G\mathcal{M})\sim\mathcal{O}(1)$ and $a_{\mu}=0.1$, consistent with the non-relativistic approximation [44, 52]. Note that $\tau_{s}$ is a separate timescale from the oscillation period of the cloud $\tau\sim\mu_{b}^{-1}$, and therefore, a quasi-stationary approximation is valid. The evolution of the superradiant instability unfolds in two stages: first, the axionic condensate evolves due to superradiance around the BH background, on a timescale given by $\tau_{s}$ (2). Then, the cloud emits energy via GWs, a process characterized by the relevant timescale $\tau_{GW}$. There is an important separation of these two scales ($\tau_{GW}\gg\tau_{s}$), ensuring two basic physical features; i) the cloud can evolve to saturation without being depleted through GWs emission and ii) the axionic-condensate is very long lived, as supported by various perturbative calculations of GW emission by the cloud [55, 56, 57, 58]. Since the lifetime $T$ of the cloud has to take into account subprocesses with widely separated time scales, it is considered to be a phenomenological parameter.
@@
where $A$ denotes the coupling constant, and $\widetilde{R}_{\mu\nu\rho\sigma}=\frac{1}{2}R_{\mu\nu\alpha\beta}\ \epsilon^{\alpha\beta}\!_{\rho\sigma}\$is the dual of the Riemann tensor, with $\epsilon_{\mu\nu\rho\sigma}$ the covariant Levi-Civita tensor.
@@
where $\hat{\alpha}_{\lambda,\vec{k}},\hat{\alpha}^{\dagger}_{\lambda,\vec{k}}$ denotes the dimensionless annihilation/creation operators obeying the usual commutation relations, $\hat{\alpha}_{\lambda,\vec{k}},\hat{\alpha}^{\dagger}_{\lambda^{\prime},\vec{k}^{\prime}}]=\delta_{\lambda\lambda^{\prime}}\delta_{\vec{k}\vec{k}^{\prime}}$, where $\lambda=L,R$ denotes left (L) and right (R) polarisation respectively, $f_{k}$ is the (dimensionless) single graviton strain, $f_{k}\equiv\kappa(2V\Omega_{k})^{-\frac{1}{2}}$, while $\Omega_{k}\equiv k$ is the frequency of GW with momentum $\vec{k}$. The polarisation tensors obey the normalization $e^{*(\lambda)}_{ij}e_{(\lambda^{\prime})}^{ij}=2\delta_{\lambda\lambda^{\prime}}$. The interaction linear in the graviton (5) is responsible for the production of graviton coherent states [65]. The remaining interactions correspond to processes including two gravitons; the first one (6), stemming from GR, describes two axions annihilating into two gravitons, while the second one (Superradiant Axionic Black-Hole Clouds as Seeds for Graviton Squeezing), associated with the anomaly, corresponds to an axion decay into two gravitons.
@@
[…]
@@
$I_{ij}(\vec{k},\vec{k}^{\prime})=\int d^{3}\vec{x}\ \partial_{i}\Psi_{2p}(\vec{x})\partial_{j}\Psi_{2p}(\vec{x})e^{-i(\vec{k}+\vec{k}^{\prime})\cdot\vec{x}}\,$, (14)
derived from the source. Since two axions are involved in the process, $E=2\mu_{b}$. The axion occupation number $N_{2p}$ acts as an enhancement parameter (in analogy to the quantum-optics SFWM). The angular and polarisation correlations of the gravitons emitted from the axionic cloud are determined from (13). Considering gravitons emitted in the yzy-z plane with $\vec{k}+\vec{k}^{\prime}$ corresponding to angles $\theta_{k+k^{\prime}}=\varphi_{k+k^{\prime}}=\pi/2$, we can express (13) in terms of the angle $\Delta\theta$ between the two gravitons (c.f. figure 2).
Figure 2: Angular and polarisation correlations for the GR interaction. The $2p$-state results in the asymmetry between the $L$ and $R$ pairs. The plot corresponds to $a_{\mu}=0.1$.
@@
$\times\left([e^{(3)}(\vec{k}^{\prime})]_{m}\ e^{(\lambda)}_{mj}(\vec{k})\ e^{(\lambda^{\prime})}_{jl}(\vec{k}^{\prime})[e^{(3)}(\vec{k})]_{l}\ +\right$.
@@
Figure 3: Angular and polarisation correlations for the CS interaction. Only pairs of opposite polarisations are produced; Maximal entanglement occurs between the L and R polarisations. The plot corresponds to $a_{\mu}=0.1$.
@@
$\times\sum_{\lambda,\lambda^{\prime}}\int d\tilde{k}d\tilde{k}^{\prime}\ \tilde{k}\tilde{k}^{\prime}\delta\left(\tilde{k}+\tilde{k}^{\prime}-2\right)\Big{|}\widetilde{\mathcal{I}}^{(GR)}_{(\lambda,\vec{k})(\lambda^{\prime},\vec{k}^{\prime})}\Big{|}^{2}\ ,$ (21)
where $\tilde{k}^{(\prime)}=k^{(\prime)}/\mu_{b}$ and $\mathcal{I}=\widetilde{\mathcal{I}}/\mu_{b}^{2}$. In order to estimate the number of gravitons in the squeezed vacuum state, we consider $\vec{q}=-(\vec{k}+\vec{k}^{\prime})$ lying on $\theta_{k+k^{\prime}}=\varphi_{k+k^{\prime}}=\pi/2$ (c.f. figure 2). The main contribution to the integral (21) comes from momentum vectors $\vec{k}$, $\vec{k}^{\prime}$ satisfying $k\approx k^{\prime}\approx\mu_{b}$, whilst the maximum contribution corresponds to their relative angle $\Delta\theta=0.964\pi$. The integrations over the solid angles will yield a factor at most of order $\left(4\pi\right)^{2}$, implying the following upper limit,
@@
Note here that the large suppression induced by the ratio $\big{(}\mu_{b}/M_{\rm Pl}\big{)}$ is compensated by the number of axions in the cloud , $N_{2p}\left(\mu_{b}/M_{\rm Pl}\right)^{2}\approx 10^{-3}a_{\mu}^{2}=10^{-5}\$(see (3)).
This is not the case for the anomaly induced interaction. In an effective field theory, the CS-coupling to the axion is suppressed below the cut-off scale of the theory. The pertinent squeezing effect is proportional to the coupling strength of the axion to the (gravitational) CS anomaly. For the string-inspired KR axion coupling [47] $A\sim 10^{-2}M_{\rm Pl}/M^{2}_{\text{s}}$, with $M_{\text{s}}$ the string scale. As in the GR case, we argue (c.f. figure 3):
@@
It is evident from (22) and (23) that the lifetime $T$ of the axionic cloud is also a parameter able to induce significant squeezing. In contrast to the short lived quasi-normal modes (QNMs) [73], the longevity of axionic clouds seems to overcome limitations and produce appreciable squeezing effects. We shall identify $T$ with the lifetime of the cloud, $T=\tau_{\text{cloud}}$. The lifetime of the axionic condensate is estimated in the literature by means of perturbative calculations and analytical approximations [55, 56, 57, 74, 44], and all seem to agree that there is a clear separation of scales, i.e. $\tau_{\text{cloud}}\gg\tau_{s}$. For example, in [74], the lifetime of the cloud results in a factor $T\sim 10^{7}\tau_{s}$, for which, on using (2) and  (22), yields,
$\sum_{I,J}|\mathcal{G}^{(GR)}_{IJ}|^{2}\lesssim 60\ .$ (24)
Hence, on account of (19) and (20), the total number of gravitons in the squeezed vacuum state $N_{gr}$ is given by an infinite sum of positive integer powers of quantities of order given typically by (24), which leads to an exponential enhancement of $N_{gr}$, resembling the $\langle N_{gr}\rangle=\sinh^{2}r$ behavior of the one-mode case [22]. From (24), we obtain an estimate of the upper bound of the squeezing parameter in our multi-mode squeezed-graviton case $r_{\rm multi-mode}^{2}\equiv\sum_{I,J}|\mathcal{G}^{(GR)}_{IJ}|^{2}\lesssim 60$, which would correspond to an upper bound on the average number of squeezed gravitons $\langle N_{gr}\rangle\lesssim\mathcal{O}(10^{7})$. We note at this point that, depending on the detailed parameters of the BH and axion masses, we obtain a range $\sum_{I,J}|\mathcal{G}^{(GR)}_{IJ}|^{2}={\mathcal{O}}(10-60)$, with the upper limit (24) being pretty robust in the non-relativistic case we consider here. In the relativistic case, where curvature effects cannot be avoided, the almost anti-collinear emission of graviton polarisation states breaks down, which might lead to additional enhancement of the squeezing. In contrast, the CS anomaly-induced squeezing (23) is still highly suppressed, compared to that induced by GR, since higher curvature interactions are subleading at low energy scales [13], determined by the axion mass, $\mu_{b}$.
Observational Prospects and Outlook. Before closing we would like to make some remarks on the experimental observability of the multi-mode-graviton squeezing parameters associated with (24). As already mentioned, the non-observation of squeezed single-mode graviton states by LIGO/Virgo interferometers [17, 32] implies [33] an upper bound on the pertinent squeezing parameter $r<41$. The fact that a GW involves a continuum of modes, prompted the authors of [33] to consider (the more realistic case of) a Gaussian profile of squeezed modes, with the peak corresponding to the characteristic frequency of the classical GW as measured by the detector in a given event. This is how these authors derived their estimate on the upper bound of the allowed region of $r$. If these results are applied at face value to our case, then from (22) one can constrain the axionic-cloud life time, and thus falsify models with too long lifetimes, such as the one leading to the upper bound in (24).
We note, however, that an extension of the (Gaussian) wave-packet approach of [33] to our multi-mode squeezed graviton states needs to be done before definite conclusions are reached on the allowed region of the pertinent squeezing parameters in our case, based on the LIGO/Virgo data. We also remark at this stage that an important feature of our suggested collective multi-mode graviton states is that they are entangled in polarisation. Because of this entanglement property, which is purely quantum in origin, the difficulties faced in the prospects of detecting single gravitons by simply looking for clicks in the detectors [11] might be avoided. Indeed, as discussed in [12], such “click”-effects are mimicked by single-mode classical GW, which does not apply to our case. A combination of methods suggested in [12], in this respect, with potential, currently unachieved, but not impossible, graviton-polarisation and correlation statistics measurements, might be a way forward. The above constitute important but challenging future research avenues which fall beyond the scope of our analysis here. We hope, nonetheless, that our work provides a motivation for experimental searches and theoretical feasibility studies of such phenomena in future interferometers. Given the current LIGO/Virgo sensitivities [17, 32, 33], such future instruments will have the sensitivity to falsify our models, via exclusion of too-long axion lifetimes.
Our analysis above has been restricted to the flat-background approximation and crucially depends on the longevity of the axionic clouds. Extending the situations to curved spacetimes in the neighborhood of the Kerr solution is a significant technical and conceptual challenge, which we did not discuss here. In early-Universe cosmology (inflation), spacetime curvature effects are known to lead to enhancement of the graviton squeezing. Whether such a situation persists in our local case is not known. Nonetheless we refer the reader to [62, 75], where back reaction effects of axions on the Kerr geometry in the context of CS gravity have been studied in conjunction with the strength of the CS coupling, which above a critical value leads to significant effects on the BH angular momentum. It would also be interesting but challenging to extend the approach to take into account axion self interactions. All such effects may provide additional modifications (perhaps enhancement) on graviton squeezing.